OpenStudy (anonymous):

okay guys please answer this for me: the length of a rectangle is 6 centimeters more than the width. the area is 216 square centimerters. what are the dementions??? oh and if you could work it out for me too thatd be great!!!!

OpenStudy (anonymous):

l=6w area=l*w 216=6w*w 216=6w^2 divide by 6 w^2=36 w=6 l=6w=6*6=36

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

lemme give you an example: A=lw (120cm2) = (3w + 2cm) (w) 120cm2 = 3w2 + 2cm w -3w2 - 2cm w + 120cm2 = 0 -1(-3w2 - 2cm w + 120cm2) = -1(0) 3w2 + 2cm w - 120cm2 = 0 (3w + 20cm) (w - 6cm) = 0 w - 6 = 0 w = 6cm (width) l = 3w + 2cm l = 3(6cm) + 2cm l =18cm + 2cm l = 20cm (length)

OpenStudy (anonymous):

what is i dont understand any thing whats name this Lesson

OpenStudy (anonymous):

i could figure it out but they put some stuff in the example that i dont know where they came from.

OpenStudy (anonymous):

it geometry or algebra... i dont know. im homeschooled and this is what i have to do.

OpenStudy (anonymous):

okey

OpenStudy (anonymous):

ive had like three people look at this and no one could get it. i have to take the test like friday.